OpenStudy (anonymous):

what is the volume of a cylinder when the radius is 3km and the height is 7km

OpenStudy (lucaz):

to find out the volume of a cylinder you calculate the area of the base and multiply by the height, area of the base is a circle, so you calculate it's area: \[\large{\pi\times radius^2=~\pi\times (3)^2=9\pi km^2}\]

OpenStudy (lucaz):

now the volume is the base times the height \[\large{9\pi~km^2\times 7~km=63\pi~km^3}\]